A survey indicated that 78% of secondary students graduate within five years of entering grade 9. This result is considered accurate with plus/minus 3%, 19 of 20 times. Estimate the sample size of this survey.
Answer by Boreal(15235)

The error at the 5% level is equal to z(0.975)*sqrt (p*(1-p))/n=0.03
square both sides; z=1.96
3.8416*0.78*0.22/n=0.0009
0.6592=0.0009n
n=732.46 or 733.
Check:
A rough guide at the 95% level is the error is 1/sqrt (n)=1/sqrt (733)=0.0369.
This guideline is accurate when z=2 and p=1-p=0.5. Nevertheless, this is a good check.